# Bifurcations of twisted solutions in a continuum limit for the Kuramoto model on nearest neighbor graphs

### v2

@@ 2. Previous Fundamental Results
$\mathbf{u}_{n}(t)=\sum^{n}_{i=1}u_{k}^{n}(t)\mathbf{1}_{I_{k}^{n}},$
where $\mathbf{1}_{I_{j}^{n}}$ represents the characteristic function of $I_{k}^{n}$, $i\in[n]$. Let $\|\cdot\|$ denote the norm in $L^{2}(I)$. In our setting as stated in Section 1, we slightly modify the arguments given in the proof of Theorem 2.3 of [12] or Theorem 3.1 of [21] to obtain the following.
@@
    If for any $\varepsilon>0$, there exist $\delta_{1}>0$ such that for $n>0$ sufficiently large, any solution $u_{i}^{n}(t)$, $i\in[n]$, to the KM (1.1) with
    $|u_{i}^{n}(0)-\bar{u}_{i}^{n}(0)|<\delta_{1},\quad i\in[n],$
@@
    $|u_{i}^{n}(t)-\bar{u}_{i}^{n}(t)|<\varepsilon,\quad i\in[n],$
    then $\bar{\mathbf{u}}(t)$ is stable. Moreover, If for any $\varepsilon>0$, there exists $\delta_{2}>0$ such that for $n>0$ sufficiently large, any solution $u_{i}^{n}(t)$, $i\in[n]$, to the KM (1.1) with
    $|u_{i}^{n}(t)-\bar{u}_{i}^{n}(t)-\theta|<\delta_{2},\quad i\in[n],$
    converges to $\bar{u}_{i}^{n}(t)$, $i\in[n]$, then $\bar{\mathbf{u}}(t)$ is asymptotically stable.
@@
    $\|\mathbf{u}_{n}(0)-\bar{\mathbf{u}}_{n}(0)\|<\delta,$
@@
    $\|\mathbf{u}_{n}(t)-\bar{\mathbf{u}}_{n}(t)\|<\varepsilon.$
@@
    $\lim_{t\to\infty}\lim_{n\to\infty}\|\mathbf{u}_{n}(t)-\bar{\mathbf{u}}_{n}(t)\|=0,$
    where $\mathbf{u}_{n}(t)$ is any solution to (1.1) such that $\mathbf{u}_{n}(0)$ is contained in the basin of attraction for $\bar{\mathbf{u}}(t)$.
@@ 3. Linear Eigenvalue Problem
where $\lambda\in\mathbb{C}$ is a constant and $\mathrm{id}$’ is the identity operator, has a solution $\phi(x)\in L^{2}(I)$ i.e., $\mathscr{L}-\lambda\,\mathrm{id}$ is surjective if $\lambda$ is not an eigenvalue given above, since $\phi(x),h(x)\in L^{2}(I)$ are expanded in Fourier series and the Fourier series of $\phi(x)$ is obtained.
@@
There exists a monotonically increasing positive sequence $\{\bar{\zeta}_{j}\}_{j=1}^{\infty}$ such that $\bar{\zeta}_{j}\in(j\pi,(j+1)\pi)$, $\psi_{1}(\zeta)$ is monotonically decreasing or increasing on $(\bar{\zeta}_{j},\bar{\zeta}_{j+1})$ and has a local minimum or maximum at $\zeta=\bar{\zeta}_{j}$, depending on whether $j$ is even or odd. Moreover,
@@
Letting $\zeta=2\pi q\kappa\neq 0$ and $\zeta_{1}=2\pi\ell\kappa\neq 0$, we rewrite the above relation as
$\frac{\sin(\zeta-\zeta_{1})}{\zeta-\zeta_{1}}=\frac{\sin(\zeta+\zeta_{1})}{\zeta+\zeta_{1}}=\frac{\sin\zeta}{\zeta}=c,$
@@
$\begin{split}&c\zeta\cos\zeta_{1}\mp\sqrt{1-c^{2}\zeta^{2}}\sin\zeta_{1}=c(\zeta-\zeta_{1}),\\ &c\zeta\cos\zeta_{1}\pm\sqrt{1-c^{2}\zeta^{2}}\sin\zeta_{1}=c(\zeta+\zeta_{1}),\end{split}$ (3.6)
where the upper or lower signs are taken simultaneously. Hence, we have $\cos\zeta_{1}=1$, i.e., $2\ell\kappa\in\mathbb{N}$. So it follows from (3.6) that
$c\zeta=c(\zeta-\zeta_{1})=c(\zeta+\zeta_{1}),$
which yields $c=0$ and $\sin\zeta=0$, i.e., $2q\kappa\in\mathbb{N}$. Thus we complete the proof. ∎
@@
Figure 1. Dependence of $\chi_{1}(l,q)$ on $\kappa$ for $l<2q$: (a) $q=1$; (b) $q=2$; (c) $q=3$; (d) $q=4$. It is plotted as the lines of which color is black for $l=1$, red for $l=2$, blue for $l=3$, orange for $l=4$, green for $l=5$, purple for $l=6$ and brown for $l=7$.
@@ 4. Bifurcations
In this section, we take $\kappa$ as a control parameter and analyze bifurcations of the twisted state solutions (1.8) in the CL (1.6). Let $\kappa_{\ell q}$ denote the value of $\kappa$ satisfying (3.2), i.e., $\chi_{1}(\kappa_{\ell q};\ell,q)=0$. It follows from the analysis of Section 3 that bifurcations occur at $\kappa=\kappa_{\ell q}$, $\ell\in\mathbb{N}$. In particular, when $\sigma=0$, the related linear operator $\mathscr{L}$ always has a zero eigenvalue, which is especially of geometric multiplicity three and very degenerate at $\kappa=\kappa_{\ell q}$ due to the translation symmetry stated at the beginning of Section 3, Henceforth we assume that $\ell=1$ for simplicity, Actually, $\kappa_{1q}$ is the smallest of $\kappa_{\ell q}$, $\ell\in\mathbb{N}$, for $q\in[4]$, as stated above. By Proposition 3.2(ii), $\kappa_{1q}=\zeta_{0}/2\pi$ for $q=1$.
@@ 4.1. Center manifold reduction
$\mu_{j}=p\chi_{1}(\kappa_{1q};j,q)\cos\sigma,\quad j\in[n]\setminus\{1\},$
$\nu_{j}=p\chi_{2}(\kappa_{1q};j,q)\sin\sigma,\quad j\in[n],$
@@ 4.2. Case of $\sigma=0$
Fix $q\in\mathbb{N}$ and suppose that $\mu_{j}=p\chi_{1}(\kappa_{1q};j,q)<0$ for any $j\neq 1$. If $\bar{\chi}_{1q}^{\prime}\beta_{0}>0$ ((resp. $\bar{\chi}_{1q}^{\prime}\beta_{0}<0)$, then a bifurcation of the one-parameter family of twisted solutions
@@
    The family $\mathscr{U}_{q}$ is stable for $\kappa<\kappa_{1q}$ and unstable for $\kappa>\kappa_{1q}$ near $\kappa=\kappa_{1q};$
@@
    There exists a stable ((resp. an unstable)) two-parameter family of modulated twisted solutions
@@
    for $\kappa>\kappa_{1q}$ ((resp. $\kappa<\kappa_{1q})$ near $\kappa=\kappa_{1q}$.
@@ 4.3. Case of $\sigma\neq 0$
Fix $q\in\mathbb{N}$ and suppose that $\mu_{j}=p\chi_{1}(\kappa_{1q};j,q)<0$ for any $j\neq 1$. If $\bar{\chi}_{1q}^{\prime}\beta_{\sigma}<0$ ((resp. $\bar{\chi}_{1q}^{\prime}\beta_{\sigma}>0$), then a bifurcation of the one-parameter family of twisted solutions
@@
    The family $\tilde{\mathscr{U}}_{q}$ of twisted solutions is stable for $\kappa<\kappa_{1q}$ and unstable for $\kappa>\kappa_{1q}$ near $\kappa=\kappa_{1q};$
@@
    There exists a stable ((resp. an unstable)) two-parameter family of oscillating twisted solutions
@@
    for $\kappa>\kappa_{1q}$ ((resp. $\kappa<\kappa_{1q})$ near $\kappa=\kappa_{1q}$, where
    $\tilde{\Omega}=\Omega+p\rho_{0}\sin\sigma\left(\frac{\bar{\chi}_{1q}^{\prime}(\kappa-\kappa_{1q})}{\beta_{\sigma}}\right)+O(\sqrt{r^{8}+\mu^{4}})$
@@
$\dot{\xi}_{0}=\tilde{\omega}p\rho_{0}\sin\sigma r^{2}+O(\sqrt{r^{8}+\mu^{4}}),$
@@ 5.2. Cases (ii) and (iii): Random graphs
In Figs. 14 and 15, $u_{k}^{n}(t)$, $k\in[n]$, are plotted as small red disks for cases (ii) and (iii), respectively, where the time $t=1000$ was chosen such that they may be regarded as the steady states, and the same values as in Figs. 12 and 13 for $n$, $\kappa$ and $u_{k}^{n}(0)$, $k\in[n]$ were used. Here $\sigma=0$ in plates (a) and (b) of both figures and $\sigma=\pi/3$ in plates (c) and (d). The most probable twisted state (1.8) for the CL (1.6) from the numerical result is also plotted as a blue line in each figure, as in Figs. 7 and 8. We observe that the KM (1.1) exhibits such twisted states as detected by Theorems 4.2 and 4.4 for the CL (1.6), although small fluctuations due to randomness are found. Oscillating twisted solutions approximately given by (4.13) were not observed in numerical simulations for random graphs of cases (ii) and (iii) whenq=2q=2, since they are very subtle even in deterministic graphs of case (i).
@@ Appendix A Eigenvalues of $\mathscr{L}$ when $q=0$
[…]

### v3

@@ 2. Previous Fundamental Results
$\mathbf{u}_{n}(t)=\sum^{n}_{k=1}u_{k}^{n}(t)\mathbf{1}_{I_{k}^{n}},$
where $\mathbf{1}_{I_{k}^{n}}$ represents the characteristic function of $I_{k}^{n}$, $k\in[n]$. Let $\|\cdot\|$ denote the norm in $L^{2}(I)$. In our setting as stated in Section 1, we slightly modify the arguments given in the proof of Theorem 2.3 of [12] or Theorem 3.1 of [21] to obtain the following.
@@
    If for any $\varepsilon>0$, there exist $\delta_{1}>0$ such that for $n>0$ sufficiently large, any solution $u_{k}^{n}(t)$, $k\in[n]$, to the KM (1.1) with
    $\min_{\theta\in\mathbb{S}^{1}}|u_{k}^{n}(0)-\bar{u}_{k}^{n}(0)-\theta|<\delta_{1},\quad k\in[n],\quad\mbox{a.s.}$
@@
    $\min_{\theta\in\mathbb{S}^{1}}|u_{k}^{n}(t)-\bar{u}_{k}^{n}(t)-\theta|<\varepsilon,\quad k\in[n],\quad\mbox{a.s.},$
    then $\bar{\mathbf{u}}(t)$ is stable. Moreover, if for any $\varepsilon>0$, there exists $\delta_{2}>0$ such that for $n>0$ sufficiently large, any solution $u_{k}^{n}(t)$, $k\in[n]$, to the KM (1.1) with
    $\min_{\theta\in\mathbb{S}^{1}}|u_{k}^{n}(t)-\bar{u}_{k}^{n}(t)-\theta|<\delta_{2},\quad k\in[n],\quad\mbox{a.s.}$
    converges to $\bar{u}_{k}^{n}(t)$, $k\in[n]$, then $\bar{\mathbf{u}}(t)$ is asymptotically stable.
@@
    $\min_{\theta\in\mathbb{S}^{1}}\|\mathbf{u}_{n}(0)-\bar{\mathbf{u}}_{n}(0)-\boldsymbol{\theta}\|<\delta\quad\mbox{a.s.},$
@@
    $\min_{\theta\in\mathbb{S}^{1}}\|\mathbf{u}_{n}(t)-\bar{\mathbf{u}}_{n}(t)-\boldsymbol{\theta}\|<\varepsilon\quad\mbox{a.s.}.$
@@
    $\lim_{t\to\infty}\lim_{n\to\infty}\min_{\theta\in\mathbb{S}^{1}}\|\mathbf{u}_{n}(t)-\bar{\mathbf{u}}_{n}(t)-\boldsymbol{\theta}\|=0\quad\mbox{a.s.},$
@@ 3. Linear Eigenvalue Problem
where $\lambda\in\mathbb{C}$ is a constant and $\mathrm{id}$’ is the identity operator, has a solution $\phi(x)\in L^{2}(I)$ i.e., $\mathscr{L}-\lambda\,\mathrm{id}$ is surjective if $\lambda$ is not an eigenvalue given above, since $\phi(x),h(x)\in L^{2}(I)$ are expanded in Fourier series, and the Fourier series of $\phi(x)$ is determined if those of $h(x)$ are given.
@@
There exists a monotonically increasing positive sequence $\{\bar{\zeta}_{j}\}_{j=1}^{\infty}$ such that $\bar{\zeta}_{j}\in(j\pi,(j+1)\pi)$, $\psi_{1}(\zeta)$ is monotonically increasing or decreasing on $(\bar{\zeta}_{j},\bar{\zeta}_{j+1})$ and has a local minimum or maximum at $\zeta=\bar{\zeta}_{j}$, depending on whether $j$ is odd or even. Moreover,
@@
Letting $\zeta=2\pi q\kappa\neq 0$ and $\zeta^{\prime}=2\pi\ell\kappa\neq 0$, we rewrite the above relation as
$\frac{\sin(\zeta-\zeta^{\prime})}{\zeta-\zeta^{\prime}}=\frac{\sin(\zeta+\zeta^{\prime})}{\zeta+\zeta^{\prime}}=\frac{\sin\zeta}{\zeta}=c,$
@@
$\begin{split}&c\zeta\cos\zeta^{\prime}\mp\sqrt{1-c^{2}\zeta^{2}}\sin\zeta^{\prime}=c(\zeta-\zeta^{\prime}),\\ &c\zeta\cos\zeta^{\prime}\pm\sqrt{1-c^{2}\zeta^{2}}\sin\zeta^{\prime}=c(\zeta+\zeta^{\prime}),\end{split}$ (3.6)
where the upper or lower signs are taken simultaneously. Hence, if $c\neq 0$, then $\cos\zeta^{\prime}=1$ so that
$c\zeta=c(\zeta-\zeta^{\prime})=c(\zeta+\zeta^{\prime}),$
which yields $c=0$. So we have $c=0$ so that by (3.6) $\sin\zeta,\sin\zeta^{\prime}=0$, i.e., $2q\kappa,2\ell\kappa\in\mathbb{N}$. Thus we complete the proof. ∎
@@
Figure 1. Dependence of $\chi_{1}(\kappa;l,q)$ on $\kappa$ for $l<2q$: (a) $q=1$; (b) $q=2$; (c) $q=3$; (d) $q=4$. It is plotted as the lines of which color is black for $l=1$, red for $l=2$, blue for $l=3$, orange for $l=4$, green for $l=5$, purple for $l=6$ and brown for $l=7$.
@@ 4. Bifurcations
In this section, we take $\kappa$ as a control parameter and analyze bifurcations of the twisted state solutions (1.8) in the CL (1.6). Let $\kappa_{\ell q}$ denote the value of $\kappa$ satisfying (3.2), i.e., $\chi_{1}(\kappa_{\ell q};\ell,q)=0$. It follows from the analysis of Section 3 that bifurcations occur at $\kappa=\kappa_{\ell q}$, $\ell\in\mathbb{N}$. In particular, when $\sigma=0$, the related linear operator $\mathscr{L}$ always has a zero eigenvalue, which is especially of geometric multiplicity three and very degenerate at $\kappa=\kappa_{\ell q}$ due to the translation symmetry stated at the beginning of Section 3. Henceforth, we assume that $\ell=1$ for simplicity, Actually, $\kappa_{1q}$ is the smallest of $\kappa_{\ell q}$, $\ell\in\mathbb{N}$, for $q\in[4]$, as stated above. By Proposition 3.2(ii), $\kappa_{1q}=\zeta_{0}/2\pi$ for $q=1$.
@@ 4.1. Center manifold reduction
$\mu_{j}=p\chi_{1}(\kappa_{1q};j,q)\cos\sigma,\quad j\in\mathbb{N}\setminus\{1\},$
$\nu_{j}=p\chi_{2}(\kappa_{1q};j,q)\sin\sigma,\quad j\in\mathbb{N},$
@@ 4.2. Case of $\sigma=0$
Fix $q\in\mathbb{N}$ and suppose that $\beta_{0}\neq 0$ and $\mu_{j}=p\chi_{1}(\kappa_{1q};j,q)<0$ for any $j\neq 1$. If $\bar{\chi}_{1q}^{\prime}>0$ ((resp. $\bar{\chi}_{1q}^{\prime}<0)$, then a bifurcation of the one-parameter family of twisted solutions
@@
    The family $\mathscr{U}_{q}$ is stable ((resp. unstable)) for $\kappa<\kappa_{1q}$ and unstable ((resp. stable)) for $\kappa>\kappa_{1q}$ near $\kappa=\kappa_{1q};$
@@
    There exists a stable or unstable two-parameter family of modulated twisted solutions
@@
    for $\kappa>\kappa_{1q}$ or $\kappa<\kappa_{1q}$ ((resp. $\kappa<\kappa_{1q}$ or $\kappa>\kappa_{1q})$ near $\kappa=\kappa_{1q}$, depending on whether $\beta_{0}>0$ or <0<0.
@@ 4.3. Case of $\sigma\neq 0$
Fix $q\in\mathbb{N}$ and suppose that $\beta_{\sigma}\neq 0$ and $\mu_{j}=p\chi_{1}(\kappa_{1q};j,q)\cos\sigma<0$ for any $j\neq 1$. If $\bar{\chi}_{1q}^{\prime}<0$ ((resp. $\bar{\chi}_{1q}^{\prime}>0$), then a bifurcation of the one-parameter family of twisted solutions
@@
    The family $\tilde{\mathscr{U}}_{q}$ is stable ((resp. unstable)) for $\kappa<\kappa_{1q}$ and unstable ((resp. stable)) for $\kappa>\kappa_{1q}$ near $\kappa=\kappa_{1q};$
@@
    There exists a stable or unstable two-parameter family of oscillating twisted solutions
@@
    for $\kappa>\kappa_{1q}$ or $\kappa<\kappa_{1q})$ ((resp. $\kappa<\kappa_{1q}$ or $\kappa>\kappa_{1q})$ near $\kappa=\kappa_{1q}$, depending on whether $\beta_{\sigma}>0$ or <0<0, where
    $\tilde{\Omega}=\Omega+p\rho_{0}\sin\sigma\left(\frac{\bar{\chi}_{1q}^{\prime}(\kappa-\kappa_{1q})}{\beta_{\sigma}}\right)+O((\kappa-\kappa_{1q})^{2})$
@@
$\dot{\xi}_{0}=\tilde{\omega}+p\rho_{0}\sin\sigma r^{2}+O(\sqrt{r^{8}+\mu^{4}}),$
@@ 5.2. Cases (ii) and (iii): Random graphs
In Figs. 14 and 15, $u_{k}^{n}(t)$, $k\in[n]$, are plotted as small red disks for cases (ii) and (iii), respectively, where the time $t=1000$ was chosen such that they may be regarded as the steady states, and the same values as in Figs. 12 and 13 for $n$, $\kappa$ and $u_{k}^{n}(0)$, $k\in[n]$ were used. Here $\sigma=0$ in plates (a) and (b) of both figures and $\sigma=\pi/3$ in plates (c) and (d). The most probable twisted state (1.8) for the CL (1.6) from the numerical result is also plotted as a blue line in each figure, as in Figs. 7 and 8. We observe that the KM (1.1) exhibits such twisted states as detected by Theorems 4.2 and 4.4 for the CL (1.6), although small fluctuations due to randomness are found. Oscillating twisted solutions approximately given by (4.13) were not observed in numerical simulations for random graphs of cases (ii) and (iii) when $q=2$, since they are very subtle even in deterministic graphs of case (i).
@@ Appendix A Eigenvalues of $\mathscr{L}$ when $q=0$
for each $\ell\in\mathbb{N}$. Thus, when $\kappa$ is changed, the zero eigenvalue is always simple and the remaining eigenvalues are negative or positive, depending on $\cos\sigma$ is positive or negative.